OpenStudy (anonymous):

Two girls have $76 between them. If the first gives the second $7 they would each have the same amount money. How much did each girl have?

OpenStudy (anonymous):

38 Dollars Each All You Do Is 76/2

OpenStudy (anonymous):

Think there's a bit more to it...

OpenStudy (anonymous):

so before the first one had$45 and the second one had $31 Right??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thnx to all of u

OpenStudy (anonymous):

45-7=38 31+7=38 76/2=38

OpenStudy (anonymous):

Dividing in half is a good way to get a sense of it, then you can adjust 7 up and 7 down... that's a decent approach.